### v2

# Bi-DCSpell: A Bi-directional Detector-Corrector Interactive Framework for Chinese Spelling Check

Abstract
Chinese Spelling Check (CSC) aims to detect and correct potentially misspelled characters in Chinese sentences. Naturally, it involves the detection and correction subtasks, which interact with each other dynamically. Such interactions are bi-directional, i.e., the detection result would help reduce the risk of over-correction and under-correction while the knowledge learnt from correction would help prevent false detection. Current CSC approaches are of two types: correction-only or single-directional detection-to-correction interactive frameworks. Nonetheless, they overlook the bi-directional interactions between detection and correction. This paper aims to fill the gap by proposing a Bi-directional Detector-Corrector framework for CSC (Bi-DCSpell). Notably, Bi-DCSpell contains separate detection and correction encoders, followed by a novel interactive learning module facilitating bi-directional feature interactions between detection and correction to improve each other’s representation learning. Extensive experimental results demonstrate a robust correction performance of Bi-DCSpell on widely used benchmarking datasets while possessing a satisfactory detection ability.
Bi-DCSpell: A Bi-directional Detector-Corrector Interactive Framework for Chinese Spelling Check
Anonymous ACL submission
1 Introduction
Chinese Spelling Check (CSC) aims to automatically detect and correct spelling errors in Chinese sentences (yu2014chinese; wang2018hybrid; huang2021phmospell). It serves as a foundation for a wide range of downstream applications in information retrieval (IR) and natural language processing (NLP), including Search Query Correction, Optical Character Recognition (OCR), and Essay Scoring. In recent years, CSC has garnered increasing attention from both academia and industry, emerging as a crucial task in the realm of Artificial Intelligence (zhang2020spelling; li2024mcrspell).
Most existing approaches to CSC can be categorized into two types of framework: (a) correction-only (C-only) and (b) detection-to-correction (D2C). The C-only does not contain a detection module. Instead, it directly adopts a specially designed corrector to perform the correction task hong2019faspell; zhang2020spelling; cheng2020spellgcn; li2021dcspell; guo-etal-2021-global. The most recent C-only models are inspired by masked language model (MLM) devlin2019bert, where each character to be corrected is predicted using contextual information. The second type of framework possesses both detection and correction modules but the interaction between them is single-directional in a way akin to a pipeline cheng2020spellgcn; zhang2021correcting; zhu2022mdcspell. That is, the detector first identifies the positions of errors, which are then used as prior knowledge for the correction module. While these two types of framework have been the dominant paradigms in the current CSC literature, they neglect the bi-directional interactions between detector and corrector. This oversight leads to issues such as the substitution of correct characters with incorrect ones (over-correction) or the failure to correct misspelled characters (under-correction).
Figure 1: Comparison of three types of interactions: ‘correction-only’, ‘detection-to-correction’ and ‘bi-directional interactions’. The wrongly/ground-truth corrected characters in the candidates are in red/blue.
Figure 1 presents an example to illustrate the differences among C-only, D2C, and bi-directional interactions. The input “{CJK}UTF8gbsn你的球机真厉害,我甘败下风。" contains two misspelled characters “{CJK}UTF8gbsn机(jī, machine)" and “{CJK}UTF8gbsn败(bài, fail)", which should be corrected as “{CJK}UTF8gbsn技(jì, skill)" and “{CJK}UTF8gbsn拜(bài, worship)", respectively. Figure 1(a) illustrates the C-only approaches that are typically underpinned by an MLM. Note that MLM tends to rectify a correct low-frequency collocation into a high-frequency one yang2022cospa. In this example, as “{CJK}UTF8gbsn败(bài, fail)" is more frequent in natural language than “{CJK}UTF8gbsn拜(bài, worship)", it is difficult for a C-only approach to correct “{CJK}UTF8gbsn败(bài, fail)" properly in case there is no prior error information from detection. In Figure 1(b), the D2C interaction first uses an error detector that determines “{CJK}UTF8gbsn败(bài, fail)" is wrong. The result is then fed into the correction module, which then successfully corrects “{CJK}UTF8gbsn败(bài, fail)" as “{CJK}UTF8gbsn拜(bài, worship)". However, the corrector still may not be able to correct the other misspelled character “{CJK}UTF8gbsn机(jī, machine)" as it is not pre-detected by the detector. This reflects that one-directional interaction (from detection to correction) can still lead to under-correction or over-correction.
To fill this gap, we propose a novel Bi-directional Detector-Corrector interactive framework for Spelling check (Bi-DCSpell). The detection and correction subtasks dynamically interact with each other, implying that knowledge acquired from one can be transferred to enhance the representation learning of the other. This bi-directional exchange of knowledge facilitates mutual improvement in performance. A concrete case is depicted in Figure 1(c). Different from the C-only and D2C approaches, Bi-DCSpell not only uses the detection result (“{CJK}UTF8gbsn败(bài, fail)" is wrong), for the corrector to revise “{CJK}UTF8gbsn败(bài, fail)" as “{CJK}UTF8gbsn拜(bài, worship)", but also feeds the correction result back to the detector as contextual information to help identify that “{CJK}UTF8gbsn机(jī, machine)" may also be wrong.
Specifically, our approach uses two independent encoders: one dedicated to detection and the other to correction, to capture the subtask-specific features from the input sequence. Subsequently, we introduce an interactive learning module, which comprises a series of bi-directional cross-attention layers to enable the dynamic interactions between the two subtasks. Finally, two task-specific classifiers are used to output the detection labels and generate the correction sequence. The performance of our method is evaluated on three widely used CSC benchmark datasets: SIGHAN13, SIGHAN14, and SIGHAN15. The results show that Bi-DCSpell surpasses interaction-free and uni-directional interactive baselines by 3.0% and 1.8% respectively in correction F1 score. Moreover, our method compares favorably against the current state-of-the-art approaches on all three datasets. A further ablation study demonstrates the effect of incorporating the bi-directional interactions in CSC. We also present a case study with concrete examples to illustrate the advantages of our model.
2 Related Work
2.1 Correction-only Approaches for CSC
Correction-only (C-only) is a kind of approaches in CSC, employing a corrector to directly correct the input sequence without error detection. Recent approaches along this line with a masked language model (MLM), and have achieved significant performance improvements hong2019faspell; wang2019confusionset; zhang2020spelling; zhang2021correcting; huang2021phmospell; zhu2022mdcspell. For instance, SpellGCN cheng2020spellgcn integrates phonological and visual similarity information into character classifiers using a graph network, which then feeds the graph representation into MLM. ReaLise xu2021read employs three distinct feature networks to capture phonetic, graphemic and semantic features, ultimately passing the fused representation through MLM. These methods focus on constructing CSC features and feeding them into an MLM-based corrector. In contrast, FASPell hong2019faspell leverages phonological and visual similarity features to construct a filtering model, selects the most suitable candidate Chinese characters from a pre-trained Language Model (PLM). To better formalize this type of CSC methods, we denote the model parameters as $\bm{\theta}_{c}$, and the correction inference degenerates to:
$P(y^{C}_{i}|\bm{x})=P(y^{C}_{i}|\bm{x},\bm{\theta}_{c})$ (1)
where, $y^{C}_{i}$ represents the corrected character at the $i$-th position in the input sequence $\bm{x}$, with $C$ denoting the correction sub-task identifier.
2.2 Detection-to-correction for CSC
Detection-to-correction (D2C) represents the second type of existing CSC approaches, which first use an error detector to detect the position information of misspelled characters and then feed the detection results to the corrector. Most approaches employ a conventional multiple-stage pipeline approach, such as Soft-masked BERT zhang2020spelling, MLM-phonetics zhang2021correcting and DR-CSC huang2023frustratingly. Unlike these pipeline D2C approaches, MDCSpell zhu2022mdcspell uses parallel detection and correction feature representation modules, and the corrector receives the detector’s hidden states, thus, the inference in correction incorporates the feature from both detection and correction. Accordingly, D2C approaches can be formalized as:
$P(y^{D}_{i}|\bm{x})$ $=P(y^{D}_{i}|\bm{x},\bm{\theta}_{d})$ (2)
$P(y^{C}_{i}|\bm{x})$ $=P(y^{C}_{i}|\bm{x},\bm{\theta}_{d},\bm{\theta}_{c})$
where $\bm{\theta}_{d}$ denotes the parameters of detection.
Similar to MDCSpell, our method also employs parallel detection and correction processes. However, a key distinction is that, our model facilitates a bi-directional interactive exchange of information between the detection and correction processes, i.e., the features learned from one subtask are utilized as prior knowledge for the other. We expect such bi-directional interactions would reinforce the representation learning for both subtasks.
3 Bi-directional Interactive Framework
Figure 2: Overview of our Bi-DCSpell framework. $\oplus$ is the matrix addition operation, and $\sigma$ denotes learnable control gate.
3.1 Problem Formulation
The Chinese Spelling Check (CSC) task aims to detect and correct the misspelled characters in a Chinese sentence. It involves two sub-tasks: detection and correction. Thus it becomes a multi-task learning problem. Formally, given an input textual sequence $\bm{x}=(w_{1},w_{2},\cdots,w_{n})$, where each Chinese character $w_{i}$ is taken from a vocabulary $\bm{V}$, a CSC model needs to detect whether each character is erroneous (i.e., misspelled), so as to output a sequence of detection labels $\bm{y}^{D}=(y^{D}_{1},y^{D}_{2},\cdots,y^{D}_{n})$ and generate the corresponding corrections $\bm{y}^{C}=(y^{C}_{1},y^{C}_{2},\cdots,y^{C}_{n})$, where $y^{D}_{i}\in\{0,1\}$, $y^{C}_{i}\in\bm{V}$.
Bi-directional Interaction In this paper, we hypothesize that the bi-directional interaction between detection and correction would benefit the performance improvement of two sub-tasks, i.e., alleviating the challenges of under-correction and over-correction present in current approaches. To achieve this, we have crafted an interactive learning module that facilitates bi-directional transfer of hidden knowledge (in term of features) between detection and correction. As shown in Figure 1(c), when the system detects whether the character in input text is incorrect, the detection module integrates feature information from correction for the detection process. Simultaneously, the correction module incorporates detection features from the detection module to make corrections. Therefore, the incorporation of bi-directional interactions in CSC can be formalized as computing the following probability for $y_{i}^{D},y_{i}^{C}$:
$P(y^{D}_{i},y^{C}_{i}|\bm{x})=P$ $(y_{i}^{D},y^{C}_{i}|\bm{x},\bm{\theta}_{d},\bm{\theta}_{c})$ (3)
3.2 Bi-DCSpell for CSC
We propose a Bi-directational Detector-Corrector interactive framework for Chinese spelling check, namely Bi-DCSpell. The framework comprises three components: an Encoder module, an Interactive Learning module, and a Classifier Layer. The overall structure is depicted in Figure 2.
3.2.1 Detection and Correction Encoders
To capture the subtask-specific features, two separate encoders are used, one for detection and the other for correction. The input text is first mapped to a sequence of embedding vectors by the embedding layer ${\rm Embedding}$. Then, Bi-DCSpell takes two separate encoders for detection and correction (${\rm Encoder}^{D}$ and ${\rm Encoder}^{C}$), formalized as:
$\bm{X}={\rm Embedding}(\bm{x})$ (4)
$\bm{H}^{D}_{r}={\rm Encoder}^{D}(\bm{X})$
$\bm{H}^{C}_{r}={\rm Encoder}^{C}(\bm{X}),$
where $\bm{H}^{D}_{r},\bm{H}^{C}_{r}$ denote detection-specific and correction-specific feature representations respectively, which are essentially the hidden states without any cross-task interaction. To facilitate the calculation in the interactive learning module, we keep $\bm{H}^{D}_{r}$ and $\bm{H}^{C}_{r}$ at the same size $R^{n\times d_{h}}$, where $d_{h}$ is the dimension of hidden state.
3.2.2 Interactive Learning
To model the bi-directional interactions between detection and correction tasks, we design an interactive learning module as shown in the middle part of Figure 2. It consists of a series of stacked bi-directional interaction layers similar to the Transformer vaswani2017attention, with each layer incorporating two task-specific cross-attention networks, two learnable control gates and a feed-forward network.
Task-specific Cross-attention Network. Firstly, two task-specific cross-attention networks use the feature knowledge learnt from one task to inform the representation learning for the other, respectively. Here, we use the same structure as vanilla attention vaswani2017attention in each cross-attention process, but the input involves two representation states (i.e., detection and correction) instead of one.
In detection cross-attention, the detection representation is updated by extracting corresponding information in correction features as follows:
$\tilde{\bm{H}}^{D}_{is}=softmax(\frac{\bm{Q}_{i}^{D}{\bm{K}_{i}^{D}}^{\mathsf{T}}}{\sqrt{d_{K}}})\bm{V}_{i}^{D}$ (5)
where, $\tilde{\bm{H}}^{D}_{is}$ denotes the hidden state of interactive detection, and $softmax$ is the $softmax$ function, and $\mathsf{T}$ denotes the transpose of matrix. The set of matrices (i.e., queries $\bm{Q}_{i}^{D}$, keys $\bm{K}_{i}^{D}$, and values $\bm{V}_{i}^{D}$) is calculated by linear transformations from $\bm{H}_{i-1}^{D}$ and $\bm{H}_{i-1}^{C}$. Specifically, $\bm{H}_{i-1}^{D}$ is used to compute queries $\bm{Q}_{i}^{D}$, and $\bm{H}_{i-1}^{C}$ is used to compute values $\bm{V}_{i}^{D}$ and keys $\bm{K}_{i}^{D}$, where all the parameters in the linear functions are trainable.
And, similar to the calculation process of $\tilde{\bm{H}}_{is}^{D}$, we take $\bm{H}_{i-1}^{C}$ and $\bm{H}_{i-1}^{D}$ to compute queries $\bm{Q}_{i}^{C}$, keys $\bm{K}_{i}^{C}$ and values $\bm{V}_{i}^{C}$, and then utilize the three matrices ($\bm{Q}_{i}^{C},\bm{K}_{i}^{C},\bm{V}_{i}^{C}$) to update the correction representation by extracting the error position information in detection features $\bm{H}_{i-1}^{D}$.
$\tilde{\bm{H}}^{C}_{is}$ $=softmax(\frac{\bm{Q}_{i}^{C}{\bm{K}_{i}^{C}}^{\mathsf{T}}}{\sqrt{d_{K}}})\bm{V}_{i}^{C}$ (6)
where $\tilde{\bm{H}}_{is}^{C}$ denotes the hidden state of interactive correction.
Learnable Control Gates Then, considering that the degree of interaction between the correction and the detection is different for each example, we employ two learnable control gates ($\alpha$ and $\beta$) to regulate the extent of interaction between one sub-task and the other:
$\alpha$ $=\sigma(W_{i}^{D}[\tilde{\bm{H}}_{is}^{D},\bm{H}_{i-1}^{D}]+b_{i}^{D})$ (7)
$\bm{H}_{is}^{D^{\prime}}=$ $LN(\alpha\odot\tilde{\bm{H}}_{is}^{D}+(1-\alpha)\odot\bm{H}_{i-1}^{D})$
$\beta$ $=\sigma(W_{i}^{C}[\tilde{\bm{H}}^{C},\bm{H}_{i-1}^{C}]+b_{i}^{C})$
$\bm{H}_{is}^{C^{\prime}}=$ $LN(\beta\odot\tilde{\bm{H}}_{is}^{C}+(1-\beta)\odot\bm{H}_{i-1}^{C})$
where, $(W_{i}^{*},b_{i}^{*})$ are learnable gate parameters and $LN$ denotes the layer normalization. $\bm{H}_{is}^{D^{\prime}}$ and $\bm{H}_{is}^{C^{\prime}}$ represent detection and correction hidden states after incorporating the correction/detection features in detection/correction representation learning.
Feed-forward Network. To merge the detection and correction feature knowledge, we designed a new feed-forward network inspired by Transformer. First, we concatenate the detection and correction hidden states and then feed it into a linear multiplication layer to learn the merged features.
$\bm{H}_{i}^{DC}$ $=\bm{H}_{is}^{D^{\prime}}\oplus\bm{H}_{is}^{C^{\prime}}$ (8)
$\bm{H}_{i}^{DC^{\prime}}=\bm{W}_{2}$ $[{\rm ReLU}(\bm{W}_{1}\bm{H}_{i}^{DC}+\bm{b}_{1})]+\bm{b}_{2}$
where $\oplus$ denotes the tensor concatenation. ${\rm ReLU}$ is the $relu$ activation function, and $(\bm{W}_{1},\bm{b}_{1})$, $(\bm{W}_{2},\bm{b}_{2})$ are two sets of parameters.
Finally, the merged hidden states $\bm{H}_{i}^{DC^{\prime}}$ are projected into a normalization function, and added to $\bm{H}^{D^{\prime}}_{is}$ and $\bm{H}^{C^{\prime}}_{is}$ to update the detection and correction representations, respectively.
$\bm{H}_{i}^{D}=LN(\bm{H}^{D^{\prime}}_{is}+{\bm{H}_{i}^{DC^{\prime}}})$ (9)
$\bm{H}_{i}^{C}=LN(\bm{H}^{C^{\prime}}_{is}+{\bm{H}_{i}^{DC^{\prime}}})$
where $\bm{H}_{i}^{D}$ and $\bm{H}_{i}^{C}$ are the updated feature representations. In the last bi-directional interaction layer $L$, the updated representations $\bm{H}_{L}^{D}$ and $\bm{H}_{L}^{C}$ are utilized as the ultimate states $\bm{H}^{D}$ and $\bm{H}^{C}$.
3.2.3 Classifier Layer
Hidden states $\bm{H}^{D}$ and $\bm{H}^{C}$ are fed into two task-specific classifiers to predict the detection label and generate the correction character for each corresponding character. Given the original input sentence $\bm{x}$ and its two representation ($\bm{H}^{D}$, $\bm{H}^{C}$), we can compute its probability distributions $p(\hat{y}_{i}^{D}|\bm{x})$ and $\hat{y}_{i}^{C}$ for $i$-th character.
$\hat{y}_{i}^{D}={\rm softmax}(\bm{W}_{D}\bm{h}^{D}_{i}+\bm{b}_{D})$ (10)
$\hat{y}_{i}^{C}={\rm softmax}(\bm{W}_{C}\bm{h}^{C}_{i}+\bm{b}_{C})$
where ($\bm{W}_{D},\bm{b}_{D}$) and ($\bm{W}_{C},\bm{b}_{C}$) are trainable parameters. Thus, the $\hat{y}_{i}^{D}$ and $\hat{y}_{i}^{C}$ denote output distributions of detection and correction, respectively.
3.2.4 Model Training
Given a set of manually labeled training data $\{(\bm{x}_{j},\bm{y}^{D}_{j},\bm{y}^{C}_{j})\}^{m}_{j=1}$ ($m$ denotes the number of training samples), we use cross-entropy loss as objective functions to train the above model. To unify the training process, a linear multi-objective optimization function is designed to guide the learning of the model to balance the detection and correction processes.
$\mathcal{L}^{D}=-\sum\nolimits_{j=1}^{m}\sum\nolimits_{i=1}^{n}y^{D}_{i}\log(\hat{y}^{D}_{i})$ (11)
$\mathcal{L}^{C}=-\sum\nolimits_{j=1}^{m}\sum\nolimits_{i=1}^{n}y^{C}_{i}\log(\hat{y}^{C}_{i})$
$\mathcal{L}=\lambda\mathcal{L}^{C}+(1-\lambda)\mathcal{L}^{D}$
where the $\lambda\in[0,1]$ is a transfer-balanced hyperparameter to guide model learning.
4 Experiments
4.1 Datasets and Pre-Processing
Extensive empirical evaluation is carried out on three widely used CSC datasets: SIGHAN13 wu2013chinese, SIGHAN14 yu2014overview and SIGHAN15 tseng2015introduction. We train Bi-DCSpell using four datasets, three of which are training data sets from SIGHAN, with approximately 10,000 data samples. The fourth dataset is an additional set of training data generated by an automatic method (wang2018hybrid) with 271,009 samples.
Consistent with the current literature xu2021read; zhu2022mdcspell, we use three test datasets from SIGHAN13, SIGHAN14 and SIGHAN15 to evaluate the trained model. Furthermore, we use the same pre-processing procedure and transform the characters in these datasets to simplified Chinese with OpenCC [^1] https://github.com/BYVoid/OpenCC.
The detailed implementation settings of Bi-DCSpell and evaluation metrics are provided in appendix A and appendix B.
Dataset Baseline Backbone Detection Correction
Prec. Rec. F1 Prec. Rec. F1
SIGHAN13 Soft-Masked BERT zhang2020spelling BERT 81.1 75.7 78.3 75.1 70.1 72.5
SpellGCN cheng2020spellgcn BERT 80.1 74.4 77.2 78.3 72.7 75.4
MLM-phonetics huang2021phmospell BERT 82.0 78.3 80.1 79.5 77.0 78.2
DCN wang-etal-2021-dynamic BERT 86.8 79.6 83.0 84.7 77.7 81.0
GAD guo-etal-2021-global BERT 85.7 79.5 82.5 84.9 78.7 81.6
MDCSpell zhu2022mdcspell BERT 89.2 78.3 83.4 87.5 76.8 81.8
DORM liang2023disentangled ChineseBERT 87.9 83.7 85.8 86.8 82.7 84.7
DR-CSC huang2023frustratingly ChineseBERT 88.5 83.7 86.0 87.7 83.0 85.3
PLM-FT (ChineseBERT) ChineseBERT 86.9 82.0 84.4 85.6 80.7 83.1
Bi-DCSpell (BERT) BERT 88.2 80.6 84.2 86.8 78.7 82.6
Bi-DCSpell (ChineseBERT) ChineseBERT 89.0 85.1 87.0 87.8 84.1 85.9
SIGHAN14 Soft-Masked BERT zhang2020spelling BERT 65.2 70.4 67.7 63.7 68.7 66.1
SpellGCN cheng2020spellgcn BERT 65.1 69.5 67.2 63.1 67.2 65.3
MLM-phonetics zhang2021correcting BERT 66.2 73.8 69.8 64.2 73.8 68.7
DCN wang-etal-2021-dynamic BERT 67.4 70.4 68.9 65.8 68.7 67.2
GAD guo-etal-2021-global BERT 66.6 71.8 69.1 65.0 70.1 67.5
MDCSpell zhu2022mdcspell BERT 70.2 68.8 69.5 69.0 67.7 68.3
DORM liang2023disentangled ChineseBERT 69.5 73.1 71.2 68.4 71.9 70.1
DR-CSC huang2023frustratingly ChineseBERT 70.2 73.3 71.7 69.3 72.3 70.7
PLM-FT (ChineseBERT) ChineseBERT 66.7 70.4 68.5 65.0 68.7 66.8
Bi-DCSpell (BERT) BERT 69.9 70.9 70.4 68.5 68.7 68.6
Bi-DCSpell (ChineseBERT) ChineseBERT 70.9 73.7 72.3 69.6 72.4 71.0
SIGHAN15 Soft-Masked BERT zhang2020spelling BERT 67.6 78.7 72.7 63.4 73.9 68.3
SpellGCN cheng2020spellgcn BERT 74.8 80.7 77.7 72.1 77.7 75.9
MLM-phonetics huang2021phmospell BERT 77.5 83.1 80.2 74.9 80.2 77.5
DCN wang-etal-2021-dynamic BERT 77.1 80.9 79.0 74.5 78.2 76.3
GAD guo-etal-2021-global BERT 75.6 80.4 77.9 73.2 77.8 75.4
MDCSpell zhu2022mdcspell BERT 80.8 80.6 80.7 78.4 78.2 78.3
CoSPA yang2022cospa ChineseBERT 79.0 82.4 80.7 76.7 80.0 78.3
DORM liang2023disentangled ChineseBERT 77.9 84.3 81.0 76.6 82.8 79.6
DR-CSC huang2023frustratingly ChineseBERT 82.9 84.8 83.8 80.3 82.3 81.3
PLM-FT (ChineseBERT) ChineseBERT 79.3 83.0 81.1 77.2 80.7 78.9
Bi-DCSpell (BERT) BERT 79.6 82.4 81.0 77.5 80.2 78.8
Bi-DCSpell (ChineseBERT) ChineseBERT 82.6 85.4 84.0 80.2 84.1 82.1
Table 1: Experimental results on SIGHAN13, SIGHAN14 and SIGHAN15 test sets. Each model includes sentence-level precision, recall, and F1 score for both detection and correction. Due to the incompatibility of character-level results in the original paper zhang2020spelling, the results for Soft-Masked BERT in the table are sourced from zhang2021correcting, maintaining consistency in training data and metrics with our method. We evaluated Bi-DCSpell using two backbones, Chinese-BERT-wwm (BERT) and ChineseBERT for a fair comparison.
4.2 Model Settings
Considering that Bi-DCSpell is essentially a fine-tuning model for CSC, a wide range of CSC methods based on fine-tuning are selected as comparison models [^2] Note that, due to the inherent disparity between a fine-tuned model and one trained through a combination of pre-training and fine-tuning, we do not choose the models relying on pre-training for comparison, such as the current state-of-the-art CSC model PTCSpell wei-etal-2023-ptcspell: PLM-FT is a fine-tuned PLM on the training data, which is indeed a standard C-only model and can be viewed as a simplified version of Bi-DCSpell without the interactive learning module and the detection encoder. Soft-Masked BERT zhang2020spelling, MLM-phonetics zhang2021correcting and DR-CSC huang2023frustratingly utilize a pipeline model, they all use the information of the previous step as the prior information of the next step. SpellGCN cheng2020spellgcn incorporates phonological and visual similarity knowledge representation into BERT by employing a specialized graph convolutional network. DCN wang-etal-2021-dynamic uses a dynamically connected network to measure the degree of dependence between any two adjacent Chinese characters. CoSPA yang2022cospa reports an alterable copy mechanism to increase the generation probability of the original input, thereby mitigating the over-correction. DORM liang2023disentangled allows the direct interaction between textual and phonetic information. MDCSpell zhu2022mdcspell utilizes multi-task learning to jointly model detection and correction, implementing a multi-task learning model with task-specific feature encoders similar to those in Bi-DCSpell. However, it neglects the knowledge interaction between detection and correction.
4.3 Main Results
The experimental results are shown in Table 1. Note that the performance of Bi-DCSpell depends on various hyper-parameters including $\alpha$ and $\beta$ for controlling the degree of interaction, and $\lambda$ for balancing detection and correction objectives in the loss function. The best performing results are reported in Table 1, while the performance of different settings of these hyper-parameters are discussed in more detail in Section 4.4.
Overall, our proposed Bi-DCSpell consistently achieves the best F1 score for both detection and correction on all three SIGHAN datasets. This result indicates the effectiveness of our method.
In the same scenario (i.e., backbone is ChineseBERT), compared with PLM-FT, which is a C-only baseline, Bi-DCSpell significantly improves the correction F1 score by 2.8%, 4.2%, and 3.0% respectively, validating the necessity of modeling the interactions between detection and correction.
Compared with BERT-based Soft-Masked BERT, MLM-phonetics, MDCSpell, CoSPA and ChineseBERT-based DR-CSC, which all adopt uni-directional interaction framework (i.e., D2C), Bi-DCSpell performs significantly better on all the datasets in both scenarios, confirming that modeling bi-directional interactions between detection and correction can further improve the CSC performance.
Compared with MLM-phonetics, CoSPA and SpellGCN with BERT backbone, which all explicitly introduce extra phonological and/or visual information into the inference process, the proposed Bi-DCSpell (BERT) does not incorporate any external information but it still achieves better F1 scores on almost all datasets. This result demonstrates the competitive performance and robustness of our proposed method.
Compared with DORM that modeled the direct interaction between textual and phonetic information in the same scenario, Bi-DCSpell (ChineseBERT) achieves better F1 scores on all datasets.
4.4 Ablation Study
In this subsection, we use SIGHAN15 test sets to examine the effects of various core components and parameters of our model, including the interactive learning module, the number of bi-directional interaction layers, the control gate factors $\alpha$ and $\beta$, and the balancing hyperparameter $\lambda$.
Method Interaction Det-F1 Cor-F1
PLM-FT no 81.1 78.9
D2C uni-directional 82.7+1.6 80.7+1.8
Bi-DCSpell bi-directional 84.0+2.9 81.9+3.0
Table 2: The results of three interaction frameworks on the SIGHAN15 test set. D2C is the Bi-DCSpell with the fixed gate $\alpha=0.0$.
Bi-directional interactive learning. Table 2 shows a result comparison between three different frameworks, including PLM-FT (C-only), D2C and Bi-DCSpell. The results illustrate that incorporating the interactions between detection and correction lead to significant performance improvements, and the modeling of bi-directional interactions brings more benefits than considering uni-directional interactions only.
Figure 3: Correction F1 scores of Bi-DCSpell with different combinations of interaction degrees ($\alpha$ and $\beta$). The deeper color indicates the higher performance.
Figure 4: The average value of $\alpha$ and $\beta$ with different bi-directional interaction layers.
The optimal degree of interaction. The Bi-DCSpell framework contains two learnable controlling gates: $\alpha$ and $\beta$. We first use a heatmap to visualize the impact of these two gate values (set them as parameters) on the correction F1 score, as depicted in Figure 3. In general, a moderate bi-directional interaction yields the highest correction F1 score, outperforming both unidirectional ($\alpha=0.0$ or $\beta=0.0$) and fully bi-directional interactions ($\alpha=1.0$ and $\beta=1.0$). Specifically, the model achieves the optimal correction F1 score (darkest color) at ($\alpha=0.4$, $\beta=0.3$), with progressively lighter colors when moving away from this point. In other words, the F1 score gradually decreases, reaching a minimum at ($\alpha=0.0$, $\beta=0.0$) where there is no interaction.
Then, we compute the mean values of $\alpha$ and $\beta$ across different interaction layers in the optimal model scenario, as illustrated in Figure 4. In both layers, both $\alpha$ and $\beta$ are less than 1, aligning with the findings in Figure 3, and the extent of information exchange from correction to detection is lower than from detection to correction.
Figure 5: The F1 scores for detection and correction vary as the number of bi-directional interaction layers increases.
case 1 Input {CJK}UTF8gbsn 团长将士兵布(bù)署(shǔ)在城外,让他们安(ān)兵不动。
C-only {CJK}UTF8gbsn 团长将士兵布置(bù zhì)在城外,让他们按(àn)兵不动。
D2C 0 0 0 0 0 0 0 0 0 0 0 0 0 0 1 0 0 0 0
{CJK}UTF8gbsn 团长将士兵布(bù)署(shǔ)在城外,让他们按(àn)兵不动。
Bi-DCSpell 0 0 0 0 0 1 0 0 0 0 0 0 0 0 1 0 0 0 0
{CJK}UTF8gbsn 团长将士兵部署(bù shǔ)在城外,让他们按(àn)兵不动。
Translation The captain deployed the soldiers outside the city and ordered them to stand still.
case 2 Input {CJK}UTF8gbsn 任何因(yīn)难(nán)都不能下(xià)倒(dǎo)有坚强意志的对(duì)员们。
C-only {CJK}UTF8gbsn 任何困难(kùn nán)都不能击败(jī bài)有坚强意志的队(duì)员们。
D2C 0 0 1 1 0 0 0 0 0 0 0 0 0 0 0 1 0 0 0
{CJK}UTF8gbsn 任何原因(yuán yīn)都不能下(xià)倒(dǎo)有坚强意志的队(duì)员们。
Bi-DCSpell 0 0 1 0 0 0 0 1 0 0 0 0 0 0 0 1 0 0 0
{CJK}UTF8gbsn 任何困难(kùn nán)都不能吓倒(xià dǎo)有坚强意志的队(duì)员们。
Translation No difficulty can intimidate team members with strong determination.
Table 3: Examples of CSC results from Bi-DCSpell, in comparison with results from C-only and D2C baselines. Red and blue are used to mark incorrect and correct characters, respectively.
The number of bi-directional interaction layers. As shown in Figure 5, the F1 scores of detection and correction are subject to some fluctuations when the number of bi-directional interaction layers changes between 1 and 8. These fluctuations generally increase when the number varies from 1 to 2, and then show a downward trend. Especially when the number is 2, the F1 scores reach a maximum value in both detection and correction. Therefore, considering the efficiency, we choose a layer number of 2 for the remaining experiments.
Figure 6: The F1 scores with respect to $\lambda$.
The balancing hyperparameter $\lambda$ in loss function. Figure 6 provides the detection, correction and hard detection [^3] To clarify, the evaluation of detection performance is based on the corrected sentences generated by the corrector. The evaluation of ‘hard detection’ is based directly on the results of the detection classifier layer. F1 scores while $\lambda$ increases. The result shows that the detection and correction F1 scores are $0.0$ when $\lambda$ is 00 (where the correction classifier is not trained), and the hard detection F1 score is $0.0$ when $\lambda$ is $1$ (where the detection classifier is not trained). Importantly, the hard detection rapidly increases when $\lambda$ increases from 00 to $0.2$. When the $\lambda$ goes from $0.8$ to $1.0$, the detection and correction performance are both decreasing. When the $\lambda=0.8$, the correction F1 score reaches a maximum. Thus, we set the $\lambda$ as $0.8$ for an optimum performance of correction.
4.5 Case Study
To further illustrate the effectiveness of Bi-DCSpell in comparison with the C-only and D2C baselines, we analyze two concrete cases in Table 3.
In the first case, the MLM-based C-only model mis-corrected “{CJK}UTF8gbsn布署(bù shǔ)" into “{CJK}UTF8gbsn布置(bù zhì)", a high-frequency collocation, resulting in over-correction. On the other hand, the D2C model did not correct “{CJK}UTF8gbsn布署(bù shǔ)", due to its failure in detection. This is an example of under-correction. In contrast, the Bi-DCSpell not only correctly detected “{CJK}UTF8gbsn布署(bù shǔ)" and “{CJK}UTF8gbsn安(ān)" as misspelled characters, but also made accurate corrections.
The second case involves three discrete misspellings (“{CJK}UTF8gbsn因难(yīn nán), 下倒(xià dǎo) and 对员(duì yuán)"). The C-only model correctly corrected two misspelled characters, but mis-corrected “{CJK}UTF8gbsn下倒(xià dǎo)” to a more common phrase “{CJK}UTF8gbsn击败(jī bài)” in this context, resulting in an over-correction. The D2C model was able to address some of the issues. However, its detection module failed to identify the wrong character “{CJK}UTF8gbsn下(xià)", and as a consequence its correction module was not able to correct it. Additionally, the detection module mis-identified the “{CJK}UTF8gbsn难( nán)” in “{CJK}UTF8gbsn因难(yīn nán)” as incorrect character, and in turn its correction module considered “{CJK}UTF8gbsn任何(rèn hé)” and “{CJK}UTF8gbsn原因(yuán yīn)” as a better match, leading to the mis-correction of “{CJK}UTF8gbsn因难(yīn nán)” to “{CJK}UTF8gbsn原因(yuán yīn)”. On the other hand, Bi-DCSpell successfully identified all misspelled characters and accurately corrected them, showing a better ability to tackle both over-correction and under-correction problems.
5 Conclusions
In this paper, we propose a novel CSC framework, namely Bi-DCSpell, which constitutes a bi-directional interactive detector-corrector framework, mutually enhancing the feature representation for the detection and correction sub-tasks. Comprehensive experiments and empirical analyses attest to the effectiveness of our approach. The explicit modeling of bi-directional interactions between sub-tasks also holds potential significance for analogous tasks, like grammatical error correction, which deserves further research in the future.
6 Limitations
6.1 Inconsistent Output
Due to the output layer of the detector-corrector framework having two results: detection labels and correction sentences, which occupy different distribution spaces, the issue of inconsistent outputs poses a new challenge. In Bi-DCSpell, two cross-attention networks are utilized to perform feature selection from the corrector to the detector and from the detector to the corrector. During model training, the Bi-DCSpell will dynamically adjust the two tasks from joint training. And we conducted a new comparison using two consistency metrics: character-level and sentence-level. Specifically, we compare PLM-FT (ChineseBERT) and Bi-DCSpell (ChineseBERT) on the SIGHAN15 dataset. The results are as Table 4:
Method D-F1 C-F1 C-l S-l
PLM-FT 81.1 78.9 95.8 41.9
Bi-DCSpell 84.0 82.1 97.7 45.7
Table 4: The consistent output statistics. The ‘D-F1’, ‘C-F1’, ‘C-l’ and ‘S-l’ denote metric ‘detection f1-score’, ‘correction f1-score’, ‘character-level consistency’ and ‘sentence-level consistency’.
From the above table, compared with PLM-FT (ChineseBERT), which is a fine-tuned PLM with two classifiers for corrector and detector, respectively, Bi-DCSpell achieves better consistency scores at both the character-level and sentence-level.
6.2 Language Limitation
In this work, we focus only on the spelling check of Chinese characters, because CSC is very different from other languages such as English. Specifically, (1) there are no delimiters between words. (2) Chinese have more than 100,000 characters, and about 3,500 are frequently used in daily life, and most characters have similar visual and/or similar pronunciations. However, we believe that the bi-directional interactions between detection and correction is also important for spelling check in English texts, which is worth an in-depth investigation in the future.
6.3 Running Efficiency
In our code implementation, we have not paid too much attention to the running efficiency of the proposed methods. More precisely, it is expensive to take about 10 hours on an A6000 GPU (48Gb GPU memory) to finish the training process. We think that the training efficiency can be improved by deploying the model training process on multiple GPUs and using data-parallel operations to increase the training batch size and shorten the training time.
Appendix A Implementation Details
We use PyTorch paszke2019pytorch to implement the proposed Bi-DCSpell the Transformers library wolf2020transformers. For scientific comparison with existing methods, two pre-trained language models, Chinese-BERT-wwm cui2021pre (abbreviated as BERT) and ChineseBERT sun2021chinesebert, are used to initialize the embedding layer and the correction encoder, each model has 12 transformer layers with 12 attention heads and outputs for each token a hidden representation of dimensionality 768. We initialize the weights of the detector encoder using the Transformer parameters of the bottom two layers in the pre-trained language model. For model training, AdamW loshchilov2018fixing is used as an optimizer with max epochs 20, the learning rate is set as 5e-5, and the batch size is set to 32. The training process takes about 10 hours on a single A6000 GPU (48Gb GPU memory).
Appendix B Evaluation Metrics
The general evaluation metrics could be divided into sentence-level and character-level metrics. Specifically, a prediction at the sentence-level is only deemed correct if all of the misspelled characters in the sentence have been detected or corrected. Compared with character-level evaluation, sentence-level evaluation is more strict and tends to yield lower scores. Following the common practice in the literature hong2019faspell; cheng2020spellgcn, we adopt the commonly used sentence-level precision, recall, and F1 score measures.